In a scene from a television show, a truck rolls down an incline and off a vertical cliff, falling into a river below. The truck starts from rest and rolls down the incline, which makes an angle of 24.0° below the horizontal, with a constant acceleration of 3.82 m/s2. After rolling down the incline a distance of 30.0 m, it reaches the edge of the cliff, which is 40.0 m above the river.

(a)

How much time (in s) does it take the truck to fall from the edge of the cliff to the landing point?

s

(b)

At the point where the truck lands in the river, how far is it horizontally from the edge of the cliff (in m)?

m
Answer #1

while an the incline, kinematic euotm From vと= 2a4e - 23.92mom) V215m/S Vxi =(1Sms) cosi = 13.70ms Vyi= (Sm/s) Sm 2=-6-10 m/s